A total of $16,000 is invested in 3 accounts earning 4%, 6%, and 8% annual interest. The combined interest for one year is $934. If the amount invested at 8% is three times the amount at 6%, find the amount invested at 4%.
Answer by mananth(16946)

investment @ 6%----------x
investment at 8%---------3x
investment @%4 ----------16000-(3x+x)=16000-4x
6%x+8%3x+4%(16000-4x)=934
multiply equation by 100
6x+8*3x+4(16000-4x)=93400
6x+24x+64000-16x=93400
14x=93400-64000
14x=429400
x=29400/14
x=2100
Investment 4% = 16000-4x
=16000-4*2100
=16000-8400
=7600